OpenStudy (pulsified333):

Assume that there are 13 board members: 8 females, and 5 males including Larry. There are 4 tasks to be assigned. Note that assigning the same people different tasks constitutes a different assignment. (1) Find the probability that both males and females are given a task. (2) Find the probability that Larry and at least one female are given tasks.

OpenStudy (pulsified333):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what do you have so far?

OpenStudy (pulsified333):

I have one of the answers right but I don't know which is right 1) 15360/17160 2) 864/17160

jimthompson5910 (jim_thompson5910):

let me think

OpenStudy (pulsified333):

ok

jimthompson5910 (jim_thompson5910):

so it says that exactly one of your answers is correct?

OpenStudy (pulsified333):

yeah I just don't know which

jimthompson5910 (jim_thompson5910):

hmm strange. I'm not getting either but I probably made a mistake somewhere. Let me double check

OpenStudy (pulsified333):

I think the first one is right though

jimthompson5910 (jim_thompson5910):

this is what I keep getting m = male f = female problem (1) 3 m, 1 f = (5 npr 3)*(8 npr 1) = 480 2 m, 2 f = (5 npr 2)*(8 npr 2) = 1120 1 m, 3 f = (5 npr 1)*(8 npr 3) = 1680 total = 480+1120+1680 = 3280 # of outcomes = 13 npr 4 = 17160 probability = 3280/17160 Problem (2) Larry + 2 m + 1 f = (4)*(4 npr 2)*(8 npr 1) = 384 Larry + 1 m + 2 f = (4)*(4 npr 1)*(8 npr 2) = 896 Larry + 0 m + 3 f = (4)*(4 npr 0)*(8 npr 3) = 1344 total = 384+896+1344 = 2624 # of outcomes = 13 npr 4 = 17160 probability = 2624/17160

OpenStudy (pulsified333):

For number 1 I got 17160-(1680+120)

jimthompson5910 (jim_thompson5910):

why did you compute it like that?

OpenStudy (pulsified333):

I followed this and I got that first answer but i don't understand how to do number 2

OpenStudy (pulsified333):

I remember now it was the second one thats wrong

jimthompson5910 (jim_thompson5910):

ah I see now, that's a better way to do #1

jimthompson5910 (jim_thompson5910):

why not determine the number of ways to have Larry + 3 males then subtract that result from the total (17160) to figure out how many ways to have Larry + at least one female

OpenStudy (pulsified333):

so how would you do that

jimthompson5910 (jim_thompson5910):

how many ways are there to assign tasks to Larry and 3 other males?

OpenStudy (pulsified333):

40

jimthompson5910 (jim_thompson5910):

how did you get 40?

OpenStudy (pulsified333):

well if larry has task A then that one thing. Then lets say he gets task B, that is different than task A

OpenStudy (pulsified333):

so isnt there 4 different tasks that larry could be assigned

jimthompson5910 (jim_thompson5910):

so he has 4 choices, yes

OpenStudy (pulsified333):

C(5,3)*4

jimthompson5910 (jim_thompson5910):

5-1 = 4 males left 4-1 = 3 slots left Compute P(4,3)

OpenStudy (pulsified333):

oh

OpenStudy (pulsified333):

24?

jimthompson5910 (jim_thompson5910):

then you multiply that by 4 4*P(4,3) = 4*24 = 96

jimthompson5910 (jim_thompson5910):

there are 96 ways to pick Larry + 3 other males

OpenStudy (pulsified333):

okay so 17160-96?

jimthompson5910 (jim_thompson5910):

hmm now that I think about it, that computes the number of ways to pick everything the opposite of "Larry + 3 other males" so larry is left out. Let me rethink

jimthompson5910 (jim_thompson5910):

did you see how I did problem 2 above? I'm guessing that answer didn't work?

OpenStudy (pulsified333):

I didn't try it yet

OpenStudy (pulsified333):

but yes I saw how you did it

jimthompson5910 (jim_thompson5910):

give it a try. It's probably wrong but I initially thought it was the correct way to do it

OpenStudy (pulsified333):

well i only have 4 attempts left so

OpenStudy (pulsified333):

look at how this person did number two for a similar but not the same problem http://www.jiskha.com/display.cgi?id=1236051521

OpenStudy (pulsified333):

@satellite73

OpenStudy (pulsified333):

I have no clue how to find the second answer

OpenStudy (pulsified333):

@dan815

OpenStudy (pulsified333):

@jim_thompson5910 i have no clue how to find the second answer

jimthompson5910 (jim_thompson5910):

same here. I'm still trying to decipher what that other poster wrote

jimthompson5910 (jim_thompson5910):

there are (13 npr 4) - (12 npr 4) = 5,280 ways to have 4 people chosen where Larry is one of the 4 people

jimthompson5910 (jim_thompson5910):

as for the other piece, hmm, I'm thinking there are (12 npr 3) - (4 npr 3) = 1,296 ways to pick at least one female

jimthompson5910 (jim_thompson5910):

so maybe we multiply the results to get 5,280*1,296 = 6,842,880 that result seems way too big though

OpenStudy (pulsified333):

it is because its bigger than the total number of ways

jimthompson5910 (jim_thompson5910):

good point

OpenStudy (pulsified333):

I found it :D 216/715

jimthompson5910 (jim_thompson5910):

I'm curious as to how you got that

OpenStudy (pulsified333):

My friend helped me with that so he's gonna explain it to me tomorrow thank you for trying though